OpenStudy (anonymous):

karen is financing 291,875 to purchase a house. she obtained a 15/5 balloon mortgage at 5.35%. what will her balloon payment be?

OpenStudy (anonymous):

please help

OpenStudy (tkhunny):

1) Calculate the monthly payment for a 15 year mortgage of $291,875 at 5.35% annual interest. 2) Let's see step one, first.

OpenStudy (anonymous):

i did ant i got 2369.40

OpenStudy (tkhunny):

Odd. Why is that a little too much? You're about $7 high.

OpenStudy (anonymous):

hmm my calculator stinks. what answer are you getting?

OpenStudy (tkhunny):

What settings did you use? Should be n = 180 pv = 291875 i = 0.0535. m = 12 fv = 0 Type = End of Period You may have no "m" and you will need to supply the monthly "I" by yourself. I = 0.0535 / 12 = 0.00445833...

OpenStudy (anonymous):

ok im confused. my choices are, 220,089.34 219112.46 246181.39 230834.98

OpenStudy (anonymous):

I chose the first one.

OpenStudy (tkhunny):

Calculate that payment. Don't give up and start staring at the answers. The right answer IS in there.

OpenStudy (anonymous):

so... i got 2.260775957. I am using the balance formula for balloon payment but its not working

OpenStudy (tkhunny):

What is 2.260775957? You need that payment. Keep at it until you get the correct: $2,361.69

OpenStudy (anonymous):

ok that isnt one of the answer choices and what kind of calculator do you have?

OpenStudy (tkhunny):

I don't have a calculator, per se. I don't normally use built in functions. I don't expect you to be able to do that, however. The payment is clearly not the answer to a problem that requires a balance. You need the correct payment, then you can do anything. 1) Calculate the monthly payment for a 15 year mortgage of $291,875 at 5.35% annual interest. $2,361.69 2) Calculate the mortgage amount for a 10 year mortgage as 5.35% annual interest, paying $2,361.69 monthly.

OpenStudy (anonymous):

ok so when i use the balloon formula im going to use 2361.69 for the m the i is going to be 0.0535 and the t is going to be 15 years?

OpenStudy (tkhunny):

?? Please state the "Balloon Formula". Let's see if we can make sense of it.

OpenStudy (anonymous):



OpenStudy (tkhunny):

That makes no sense at all. Annual Accumulation of the original Sum: \(291875\cdot (1.0535)^{5}\) Monthly Accumulation of Payments: \(2361.69\cdot\dfrac{\left(1 + \dfrac{0.0535}{12}\right)^{60} - 1}{\left(\dfrac{0.0535}{12}\right)}\) Now subtract. You DO have to know what all the parts mean.

OpenStudy (anonymous):

ok i got 160205.83. now what would i subtract?

OpenStudy (tkhunny):

For which piece? Annual accumulation of value or monthly accumulation of payments?

OpenStudy (anonymous):

monthly accumalation for anual i got 378764.7878. then i would subtract the 160205.83. which gives me 21558.9578

OpenStudy (tkhunny):

Except that 160205.83 is incorrect. Try that one again.

OpenStudy (tkhunny):

The first piece is also wrong. This interest is compounded monthly. Why are we accumulating the value annually? \(291875\cdot\left(1 + \dfrac{0.0535}{12}\right)^{60} = 381164.519935256\)

OpenStudy (anonymous):

so how is this answer correct? It's not one of my choices and just doesn't make sense to me... @tkhunny @50shadesofchocolate